OpenStudy (anonymous):

Express the percent as a fraction or mixed number in simplest form. 0.1%

OpenStudy (anonymous):

Well, that means 0.1 out of a 100. This cannot be done in fractional form, so it would be 1/1000

OpenStudy (mayankdevnani):

@bowbow always remember: \[\huge 0.1 percent=\frac{0.1}{100}\] and solve it??

OpenStudy (mayankdevnani):

can you solve it??? @bowbow

OpenStudy (anonymous):

mhm i got it

OpenStudy (mayankdevnani):

what is the answer?

hero (hero):

\[0.1\% = \frac{1}{10} \times \frac{1}{100}\]